OpenStudy (destinykiara99):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = -16t2 + 576t. After how many seconds does the projectile take to reach its maximum height? (Show your work)

OpenStudy (destinykiara99):

@myininaya would you mind helping with one last one??

myininaya (myininaya):

it is asking you to find the vertex

myininaya (myininaya):

actually it is asking you only for the x-coordinate (t-coordinate in this case) of the vertex

myininaya (myininaya):

do you know how to write it in vertex form or do you know the vertex formula

myininaya (myininaya):

if not that is okay we can work together to get there

OpenStudy (destinykiara99):

Yeah, I'll need some help because I'm not sure

myininaya (myininaya):

the vertex form of the parabola is \[h(x)=a(x-h)^2+k \text{ where the vertex is } (h,k)\] this if the form we want it in

myininaya (myininaya):

first step factor the number that is in front of the squared variable out of the term with the squared variable and the variable that is we are going to do this first: \[h(t)=-16t^2+576t \\ h(t)=-16(t^2+\frac{576}{-16}t)\] this is only the first step I didn't simplify -576/16 yet because I couldn't do it in my head

myininaya (myininaya):

that should be -36 \[h(t)=-16t^2+576t \\ h(t)=-16(t^2-36t)\]

myininaya (myininaya):

now let's just look at t^2-36t....

myininaya (myininaya):

this is the part we want to complete the square on

myininaya (myininaya):

what is k here... (go back to the first post we were in to determine this)

myininaya (myininaya):

this thing and here are some more examples these examples are when the k is negative (like for our question) \[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \\ x^2-2x+1=(x-1)^2 \\ x^2-4x+4=(x-2)^2 \\ x^2-6x+9=(x-3)^2 \\ x^2-8x+16=(x-4)^2 \\ x^2-10x+25=(x-5)^2 \\ x^2-12x+36=(x-6)^2 \\ x^2-14x+49=(x-7)^2\]

myininaya (myininaya):

it doesn't matter t or x whatever

myininaya (myininaya):

\[t^2-36t+... \text{ compare this to } t^2+kt+...\] what does k need to be

OpenStudy (destinykiara99):

36 ?

myininaya (myininaya):

-36

OpenStudy (destinykiara99):

oh, yeah, sorry

myininaya (myininaya):

so this is good you determine what k needs to be input that k into this formula: \[t^2+kt+(\frac{k}{2})^2=(t+\frac{k}{2})^2 \] everywhere you see k replace it with -36

myininaya (myininaya):

\[t^2-36t+(\frac{-36}{2})^2=(t+\frac{-36}{2})^2\] this what that should look like before simplifying

myininaya (myininaya):

can you simplify

OpenStudy (destinykiara99):

(t-18)^2

myininaya (myininaya):

so you have \[t^2-36t+(-18)^2=(t-18)^2 \\ \text{ or } t^2-36t+18^2=(t-18)^2 \\ \text{ or } t^2-36t+324=(t-18)^2\]

OpenStudy (destinykiara99):

the first one?

myininaya (myininaya):

oh no (that wasn't a multiple choice question just now... those 3 equations are all equivlanent... sorry ) ok this is going to help us with our problem let me go up and fetch it... this is what we left off with \[h(t)=-16t^2+576t \\ h(t)=-16(t^2-36t) \] anyways let's use the last simplified equation to help us complete the square here inside the parenthesis you see we need to add 324 inside the parenthesis but this means we also need to subtract 324 inside the parenthesis like this \[h(t)=-16(t^2-36t+324-324) \\ \text{ now remember the distributive property } a(b+c)=ab+ac \\ \text{ I'm going \to do that here } \\ h(t)=-16(t^2-36t+324)-16(-324)\] do you know why I did that? or do you see what follows from doing this?

myininaya (myininaya):

if you need a hint let me know

OpenStudy (destinykiara99):

Just confused on where the 324 comes from

myininaya (myininaya):

do you remember \[t^2-36t+(-18)^2=(t-18)^2 \\ \text{ or } t^2-36t+18^2=(t-18)^2 \\ \text{ or } t^2-36t+324=(t-18)^2\]

myininaya (myininaya):

18^2 is 324 this is why we did the whole 324-324 into the parenthesis so we can complete the square

myininaya (myininaya):

\[t^2+kt+(\frac{k}{2})^2=(t+\frac{k}{2})^2 \\ \text{ we had determined our } k \\ \text{ was } -36 \\ \text{ so we know the following equation is true } \\ t^2-36t+(\frac{-36}{2})^2=(t+\frac{-36}{2})^2 \]

myininaya (myininaya):

-36/2 is -18 and (-18)^2 is the same as 18^2 which is the same as 324

myininaya (myininaya):

so this is how we go this: \[t^2-36t+324=(t-18)^2 \] and that is where the 324 comes into play into our function

myininaya (myininaya):

but whatever we add into our function we must also take out

myininaya (myininaya):

\[h(t)=-16t^2+576t \\ h(t)=-16(t^2-36t) \\ h(t)=-16(t^2-36t+324-324) \\ \text{ we needed the 324 to complete the square } \\ \text{ so I just threw in +324 }\\ \text{ but also had to keep the function the same by putting in -324} \]

OpenStudy (destinykiara99):

Okay, I get it up to this point now

myininaya (myininaya):

\[h(t)=-16(\color{red}{t^2-36t+324}-\color{blue}{324}) \\ \text{ now the distributive property} \\ a(\color{red}{b}+\color{blue}c)=a \color{red}{b}+a \color{blue}{c} \\ h(t)=-16(\color{red}{t^2-36t+324})+-16(\color{blue}{-324})\]

myininaya (myininaya):

guess what we can replace the t^2-36t+324 with ?

myininaya (myininaya):

the whole point of adding the 324 was to complete the square remember that we said t^2-36t+324 is the same as .... ( instead me saying the same as you might have seen the = which means the same thing )

OpenStudy (destinykiara99):

(t-18)^2

myininaya (myininaya):

yep yep yep!! :)

myininaya (myininaya):

\[h(t)=-16(t-18)^2+-16(-324) \\ \]

myininaya (myininaya):

we really don't need to simplify further to find the x-coordinate of the vertex

myininaya (myininaya):

but if you want we can

myininaya (myininaya):

-16(-324) is...

OpenStudy (destinykiara99):

5,184

myininaya (myininaya):

\[h(t)=-16(t-18)^2+5184 \\ \text{ compare this to } \\ h(t)=a(t-h)^2+k\] the vertex is (h,k)

myininaya (myininaya):

h will be your answer

myininaya (myininaya):

whatever h is :)

OpenStudy (destinykiara99):

18! so it takes 18 seconds for it to reach maximum height!

myininaya (myininaya):

right on... and the max height is 5184 (I know the question did not ask this)

OpenStudy (destinykiara99):

Thank you so much

myininaya (myininaya):

actually 5184 feet but whatever

myininaya (myininaya):

no problem

myininaya (myininaya):

I know that was really long so give yourself a pat on the back for hanging in there that long most people would give up that should make your feel pretty successful right now anyways peace